OpenStudy (anonymous):

A person is considered to have certain symptoms when his or her temp is higher than 100.04 degrees Fahrenheit the formula F=9/5C+32 can be used to convert Celsius temps to Fahrenheit temps. For which Celsius temps is the person considered to have the symptom.

OpenStudy (anonymous):

F=9/5C+32 100.04=9/5C+32 100.04-32=9/5C (68.04=9/5C)5 340.2=9C divide both sides by 9 37.8=C

OpenStudy (anonymous):

You are awesome!

OpenStudy (anonymous):

Thanks, glad I could help(: